OpenStudy (anonymous):

Solve with matrices. 5x -3y = 18 and 2x + 7y = -1

OpenStudy (dumbcow):

elimination would be better

OpenStudy (anonymous):

|5 -3| 18| |2 7 | -1|

OpenStudy (anonymous):

row operations will give you: | 1 0| 3 | | 0 1|-1| do you see your answer?

OpenStudy (anonymous):

No, lol

OpenStudy (anonymous):

it's means: x=3 and y=-1 OK now?

OpenStudy (anonymous):

|5 -3| 18| = |10 -15 |26 | = | 5 -3 | 18 | = |5 0 | 15| |2 7 | -1| |-10 -35| 5 | |0 -41| 41| |0 1 | -1| and so on

OpenStudy (anonymous):

|5 0 | 15| is the final answer? |0 1 | -1|

OpenStudy (anonymous):

you simplify this one to: | 1 0| 3 | | 0 1|-1| and final answer in x & y: x=3 (top row) and y=-1 (lower row)